March 17, 2000

# 5 Fifth Ave. New Font City, New Fontiac United States of Fontagraphy 12345-6789

Pyotr Ilyitch Tchyfontski Research Associate Professor Division of Fontal Conversion St. Leninberg University of the New State

Tovarich Pyotr:

A line that will differ from case to case. In terms of the relationships between font descriptions and musical typography, it seems to me that only in the case of the early 1800s can even the slightest case be made for a relationship.

Yours in fontal health,

Ludwig von Beetfonten Research Associate Professor Division of Font Music Insertions Vienna University # 5 Fifth Ave. New Font City, New Fontiac United States of Fontagraphy 12345-6789

Carmen “Cacciatore” Rodriguez Full Assistant Associate Professor Division of Chicken Informatics Cuba National University Guantanemo Bay, Cuba

Senora Rodriguez:

Of course, it doesn’t have to, but it will anyway. In terms of the relationships between font descriptions and musical typography, it seems to me that only in the case of the early 1800s can even the slightest case be made for a relationship.

Yours in fontal health,

Ludwig von Beetfonten Research Associate Professor Division of Font Music Insertions Vienna University